OpenStudy (anonymous):

At noon, ship A is 20 nautical miles due west of ship B. Ship A is sailing west at 23 knots and ship B is sailing north at 21 knots. How fast (in knots) is the distance between the ships changing at 7 PM? The distance is changing at _____________ knots. (Note: 1 knot is a speed of 1 nautical mile per hour.)

OpenStudy (anonymous):

hey @pitamar do you how to solve this problem?

OpenStudy (anonymous):

Ok, let's try =) Seems like what we need here is phytagorean theorem and derivatives of polynomials. are you familiar with them both?

OpenStudy (anonymous):

i think so. pythagorean theorem is x^2 +y^2 = s^2 and to find the derivatives you do something like 2sds/dt = 2xdx/dt + 2ydy/dt. is that right?

OpenStudy (anonymous):

sorry was cooking heh well, let's start with the pythagorean theorem first. We have:  right?

OpenStudy (anonymous):

ok yes

OpenStudy (anonymous):

Ok, so we can say using pythagorean theorem that: $$ (\text{distance})^2 = (20 + 23t)^2 + (21t)^2 $$ Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wanna multiply it and see what you get? =)

OpenStudy (anonymous):

d^2 = 400 + 920t +529t^2

OpenStudy (anonymous):

ehmm, why 529? We have \((23t)^2 + (21t^2) = t^2 \cdot ( (23)^2 + (21)^2) = 970t^2\) The whole thing is pretty much $$ (\text{distance})^2 = 970t^2 + 920t + 400 $$

OpenStudy (anonymous):

So, we can take root and have $$ \text{distance} = \sqrt{970t^2 + 920t + 400} $$Because distance can only be positive ok?

OpenStudy (anonymous):

ohhh i forgot to add the 21t^2. i just foiled the first part

OpenStudy (anonymous):

lol that's ok =) So now let's call distance 'D' and we want to find the rate of change of the distance 'D' with respect to time 't'. Or dD/dt. Which is the derivative of what we have here with respect to t $$ \frac{dD}{dt} = \left( \sqrt{970t^2 + 920t + 400} \right)' = ? $$

OpenStudy (anonymous):

1/2 (1/(sqrt(970t^2 + 920t + 400)) * 1940t+920

OpenStudy (anonymous):

ye seems ok. now we want to find that in 7PM. we said t0 is at noon, so 7pm is 7 hours later where t=7

OpenStudy (anonymous):

just plug in? 1/2 (1/(sqrt(970(7)^2 + 920(7) + 400)) * 1940(7)+920

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

i got 31.09269309

OpenStudy (anonymous):

ye me too

OpenStudy (anonymous):

ok yay thank you!!

OpenStudy (anonymous):

sure, hope it's right =)

OpenStudy (anonymous):

it is :)